# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@ 14.1 Application of the principle of typicality
In the Wigner-Deutsch collaboration, the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by (77) in our world $\omega$. This is because the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ while the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. Consequently, the projector $P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)$ corresponds to (83) in this case. In contrast, the projector $P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)$ corresponds to (84) in this case. Let us calculate $P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$. Then using (76) we have that
@@
for every $n\in\mathbb{N}^{+}$. However, it follows from (74) that $\alpha(n)\neq\beta(n)$ for half of $n\in\mathbb{N}^{+}$, since
@@
Therefore, the vector (85) becomes 0 for such infinitely many $n$. Hence, the vector (85) does not become the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, since the state vector must not be a zero-vector. In contrast, the vector $P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$, which the original Postulate 7 requires to be the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, is non-zero for all $n\in\mathbb{N}^{+}$, since it equals (78) for all $n\in\mathbb{N}^{+}$. This observation supports the validity of Postulate 7. ∎
@@
Let us consider the validity of Postulate 7 further. Postulate 7 can be applied on the premise that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$. We will demonstrate the necessity of this premise, based on the Wigner-Deutsch collaboration. We call a variant of Postulate 7 where this premise is eliminated the naive Postulate 7. We will show that the naive Postulate 7 is too strong and leads to a contradiction.
On the one hand, using the naive Postulate 6 we can derive the result (82) from (73) in the same manner as we derived the result (82) from (73) above using Postulate 7. That is, we have that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$, in the form corresponding to (9). It is then easy to see that the finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ forms measurement operators.
Now, we can use the naive Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$, under the situation that the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ is regarded as the single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments where the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ is regarded as the single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$ results in a state $|\Psi^{\mathcal{F}(\mathcal{W}+\mathcal{D})}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (56) and (55), respectively. On the other hand, it is easy to check that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{W}+\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from the naive Postulate 7 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
In contrast, we can avoid this contradiction for Postulate 7, since the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ do not form a PVM and therefore the above argument does not work for Postulate 7. Hence, in Postulate 7, the premise that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$ is necessary. ∎
@@
In this case, according to (56), the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$. The succession of the measurement $\mathcal{M}^{\mathcal{W}}$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ performed over the composite system $\mathcal{S}+\mathcal{F}$ in this state forms the repeated once of the infinite repetition of measurement. Thus, for applying the principle of typicality, we have to implement the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ by unitary time-evolution. This is already done in $U_{\mathrm{II}}$ of (61) and $U_{\mathrm{III}}$ of (62).
@@
for each $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$. We denote the set $\{0,1,2\}\times\{+,-\}$ by $\Theta$, and we define a finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
@@
$\sum_{(l,m)\in\Theta}{M^{\mathcal{W}+\mathcal{D}}_{l,m}}^{\dagger}M^{\mathcal{W}+\mathcal{D}}_{l,m}$ $=\sum_{l=0,1,2}(I_{\mathcal{S}}\otimes{M^{\mathcal{W}}_{l}}^{\dagger})\left[\sum_{m=+,-}{M^{\mathcal{D}}_{m}}^{\dagger}M^{\mathcal{D}}_{m}\right](I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})$
$=I_{\mathcal{S}}\otimes\sum_{l=0,1,2}{M^{\mathcal{W}}_{l}}^{\dagger}M^{\mathcal{W}}_{l}=I_{\mathcal{S}+\mathcal{F}}.$
Thus, the finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ forms measurement operators. Then, using (89) we see that, as a whole, the sequential applications of $U_{\mathrm{II}}$ and $U_{\mathrm{III}}$ to the state $|\Psi\rangle\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$ of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ result in the state:
@@
In this case, according to (56), the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$. Thus, the unitary operator $U_{\mathrm{whole}}$ given by (91) applying to the initial state
@@
describes the repeated once of the infinite repetition of the measurements (experiments) where the succession of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ in this order is infinitely repeated. It follows from (91) that the application of $U_{\mathrm{whole}}$ forms a single measurement, which is described by the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ and whose all possible outcomes form the set $\Theta$.
@@
$\begin{split}&\sum_{(l,m)\in\Theta}(M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle\\ &=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}))((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle)\\ &=\sum_{(k,l,m)\in\Omega}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi[k,l,m]\rangle\\ &=\sum_{(l,m)\in\Theta}\left\{\sum_{k=0,1}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\right\}\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle.\end{split}$
@@
$M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))=\sum_{k=0,1}(M_{k,l,m}|\Psi\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle$ (92)
@@
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiments) in the above setting. This $\omega$ is an infinite sequence over $\Theta$ consisting of records in the observers $\mathcal{W}$ and $\mathcal{D}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ in the above setting. Since the Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $R$-random.
We use $\beta$ and $\gamma$ to denote the infinite sequences over $\{0,1,2\}$ and $\{+,-\}$, respectively, such that $(\beta(n),\gamma(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (93) and Corollary 4.4 that
@@
for every $n\in\mathbb{N}^{+}$, i.e., $\beta$ is an infinite binary sequence. It follows from Postulate 5 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
up to the normalization factor, where the equality follows from (94). Since $\omega$ is Martin-Löf $R$-random, it follows from (95) and Theorem 4.2 that for every $(l,m)\in\Theta_{c}$ it holds that
@@
where $N_{l,m}(\sigma)$ denotes the number of the occurrences of $(l,m)$ in $\sigma$ for every $(l,m)\in\Theta_{c}$ and every $\sigma\in\Theta_{c}^{*}$. Thus, in our world $\omega$, the following holds for each $(l,m)\in\Theta_{c}$: In one-fourth of the infinite repetitions of the experiment, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle$ and the observers $\mathcal{W}$ and $\mathcal{D}$ record the values $l$ and $m$, respectively, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$. We can see that the this situation coincides with that of both Case 1 and Case 2 in Section 13 regarding the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$, although the statements of both Case 1 and Case 2 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.
Let us determine the states of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ and therefore the final states of the observer $\mathcal{W}$ is confirmed before the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Using (97), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is $|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is

### v2

@@ 14.1 Application of the principle of typicality
In the Wigner-Deutsch collaboration, the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is given by (79) in our world $\omega$. This is because the final states of the observer $\mathcal{W}$ is unchanged after the measurement by the observer $\mathcal{D}$ while the final states of the observer $\mathcal{F}$ is not unchanged after the measurement by the observer $\mathcal{D}$, as we saw above. Consequently, the projector $P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)$ corresponds to (85) in this case. In contrast, the projector $P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)$ corresponds to (86) in this case.
Let us calculate $P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle,|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$. Then using (78) we have that
@@
for every $n\in\mathbb{N}^{+}$. However, it follows from (76) that $\alpha(n)\neq\beta(n)$ for half of $n\in\mathbb{N}^{+}$. This is because using (76) we have that
@@
Therefore, the vector (87) becomes 0 for such infinitely many $n$. Hence, the vector (87) does not become the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ for such infinitely many $n$ in our world $\omega$, since the state vector must not be a zero-vector. In contrast, the vector $P\left(|\Phi^{\mathcal{W}}[\beta(n)]\rangle\right)|\Psi^{\mathcal{W}}_{\mathrm{Total}}\rangle$, which the original Postulate 7 requires to be the state of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$, is non-zero for all $n\in\mathbb{N}^{+}$, since it equals (80) for all $n\in\mathbb{N}^{+}$. This observation supports the validity of Postulate 7. ∎
@@
Let us consider the validity of Postulate 7 further. Postulate 7 can be applied on the premise that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{i-1}$ for every $i=1,\dots,n$. We will demonstrate the necessity of this premise, based on the Wigner-Deutsch collaboration. We call a variant of Postulate 7 where this premise is eliminated the naive Postulate 7. We will show that the naive Postulate 7 is too strong and leads to a contradiction.
On the one hand, using the naive Postulate 7 we can derive the result (84) from (75) in the same manner as we derived the result (84) from (75) above using Postulate 7. That is, we have that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$, in the form corresponding to (9) with $n=k=2$. It is then easy to see that the finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ satisfies the completeness equation, and therefore forms measurement operators (see (94) in Section 15 below). Moreover, we can check that the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ also satisfy the condition (10) with $n=k=2$.
However, the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ do not form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$. To see this, we first note that
$M^{\mathcal{W}+\mathcal{D}}_{l,+}=\frac{1}{\sqrt{2}}|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\langle l|\otimes\langle\Phi^{\mathcal{F}}[l]|$ (89)
holds for each $l=0,1$, which follows immediately from (60), (59), and (58). Assume contrarily that the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$. Then $M^{\mathcal{W}+\mathcal{D}}_{0,+}$, in particular, is a projector on $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$. Since $M^{\mathcal{W}+\mathcal{D}}_{0,+}\neq 0$ due to (89), there is a normalized vector $|\psi_{0}\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ such that $M^{\mathcal{W}+\mathcal{D}}_{0,+}|\psi_{0}\rangle=|\psi_{0}\rangle$. Thus, using (89) we have that
$1=\langle\psi_{0}|M^{\mathcal{W}+\mathcal{D}}_{0,+}|\psi_{0}\rangle=\frac{1}{\sqrt{2}}\langle\psi_{0}|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle(\langle 0|\otimes\langle\Phi^{\mathcal{F}}[0]|)|\psi_{0}\rangle.$
It follows that $1\leq 1/\sqrt{2}$, since $\left\lvert\langle\psi_{0}|\Psi^{\mathcal{S}+\mathcal{F}}[+]\rangle\right\rvert,\left\lvert(\langle 0|\otimes\langle\Phi^{\mathcal{F}}[0]|)|\psi_{0}\rangle\right\rvert\leq 1$. Hence, we have a contradiction, and therefore the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ cannot form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$.
Now, we can use the naive Postulate 7 to determine the whole states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$, under the situation that the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ is regarded as the single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{D}}_{\mathrm{init}}\rangle$ in the repeated once of the experiments where the combination of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ is regarded as the single measurement $\mathcal{M}^{\mathcal{W}+\mathcal{D}}$ results in a state $|\Psi^{\mathcal{F}(\mathcal{W}+\mathcal{D})}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (57) and (56), respectively. On the other hand, it is easy to check that, according to Definition 7.3, the final states of the observer $\mathcal{F}$ are not unchanged after the measurement by the observer $\mathcal{W}+\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Thus, it follows from the naive Postulate 7 that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
In contrast, we can avoid this contradiction for Postulate 7, since the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ do not form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ and therefore the above argument does not work for Postulate 7. Hence, in Postulate 7, the premise that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{i-1}$ for every $i=1,\dots,n$ is necessary. ∎
@@
In this case, according to (57), the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$. The succession of the measurement $\mathcal{M}^{\mathcal{W}}$ and then the measurement $\mathcal{M}^{\mathcal{D}}$ performed over the composite system $\mathcal{S}+\mathcal{F}$ in this state forms the repeated once of the infinite repetition of measurements. Thus, for applying the principle of typicality, we have to implement the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ by unitary time-evolution. This is already done in the form (63) of $U_{\mathrm{II}}$ and the form (64) of $U_{\mathrm{III}}$.
@@
for each $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$.
We denote the set $\{0,1,2\}\times\{+,-\}$ by $\Theta$, and we define a finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ of operators acting on the state space $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{\mathcal{F}}$ of the composite system $\mathcal{S}+\mathcal{F}$ by
@@
$\begin{split}\sum_{(l,m)\in\Theta}{M^{\mathcal{W}+\mathcal{D}}_{l,m}}^{\dagger}M^{\mathcal{W}+\mathcal{D}}_{l,m}&=\sum_{l=0,1,2}(I_{\mathcal{S}}\otimes{M^{\mathcal{W}}_{l}}^{\dagger})\left[\sum_{m=+,-}{M^{\mathcal{D}}_{m}}^{\dagger}M^{\mathcal{D}}_{m}\right](I_{\mathcal{S}}\otimes M^{\mathcal{W}}_{l})\\ &=I_{\mathcal{S}}\otimes\sum_{l=0,1,2}{M^{\mathcal{W}}_{l}}^{\dagger}M^{\mathcal{W}}_{l}=I_{\mathcal{S}+\mathcal{F}}.\end{split}$ (94)
Thus, the finite collection $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ satisfies the completeness equation, and therefore forms measurement operators. Then, using (92) we see that, as a whole, the sequential applications of $U_{\mathrm{II}}$ and $U_{\mathrm{III}}$ to the state $|\Psi\rangle\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle$ of the composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$ result in the state:
@@
In this case, the state of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the ‘measurement’ $\mathcal{M}^{\mathcal{F}}$ is given by $U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)$, as we stated above. Thus, the unitary operator $U_{\mathrm{whole}}$ given by (95) applying to the initial state
@@
describes the repeated once of the infinite repetition of the measurements (experiment) where the succession of the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ in this order is infinitely repeated. It follows from (95) that the application of $U_{\mathrm{whole}}$ forms a single measurement, which is described by the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ and whose all possible outcomes form the set $\Theta$.
@@
$\begin{split}&\sum_{(l,m)\in\Theta}(M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle)))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle\\ &=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}}))((U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))\otimes|\Phi^{\mathcal{W}+\mathcal{D}}_{\mathrm{init}}\rangle)\\ &=(U_{\mathrm{III}}\circ(I_{\mathcal{S}}\otimes U_{\mathrm{II}}\otimes I_{\mathcal{D}})\circ(U_{\mathrm{I}}\otimes I_{\mathcal{W}}\otimes I_{\mathcal{D}}))(|+\rangle\otimes|\Phi_{\mathrm{init}}\rangle)\\ &=\sum_{(k,l,m)\in\Omega}(M_{k,l,m}|+\rangle)\otimes|\Phi[k,l,m]\rangle\\ &=\sum_{(l,m)\in\Theta}\left\{\sum_{k=0,1}(M_{k,l,m}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle\right\}\otimes|\Phi^{\mathcal{W}+\mathcal{D}}[l,m]\rangle.\end{split}$
@@
$M^{\mathcal{W}+\mathcal{D}}_{l,m}(U_{\mathrm{I}}(|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle))=\sum_{k=0,1}(M_{k,l,m}|+\rangle)\otimes|\Phi^{\mathcal{F}}[k]\rangle$ (96)
@@
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurements (experiment) in the above setting. This $\omega$ is an infinite sequence over $\Theta$ consisting of records in the observers $\mathcal{W}$ and $\mathcal{D}$ which is being generated by the infinite repetition of the measurement described by the measurement operators $\{M^{\mathcal{W}+\mathcal{D}}_{l,m}\}_{(l,m)\in\Theta}$ in the above setting. Since the Bernoulli measure $\lambda_{R}$ on $\Theta^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $R$-random.
We use $\beta$ and $\gamma$ to denote the infinite sequences over $\{0,1,2\}$ and $\{+,-\}$, respectively, such that $(\beta(n),\gamma(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then, since $\omega$ is Martin-Löf $R$-random, it follows from (97) and Corollary 4.4 that
@@
for every $n\in\mathbb{N}^{+}$, i.e., $\beta$ is an infinite binary sequence. Thus, it follows from Postulate 5 and (95) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$, i.e., the $n$th copy of a composite system consisting of the system $\mathcal{S}$ and the three observers $\mathcal{F}$, $\mathcal{W}$, and $\mathcal{D}$, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is given by
@@
up to the normalization factor, where the equality follows from (98). Since $\omega$ is Martin-Löf $R$-random, it follows from Theorem 4.2 and (99) that for every $(l,m)\in\Theta_{c}$ it holds that
@@
where $N_{l,m}(\omega\!\!\restriction_{n})$ denotes the number of the occurrences of $(l,m)$ in the prefix of $\omega$ of length $n$. Thus, in our world $\omega$, the following holds for each $(l,m)\in\Theta_{c}$: In one-fourth of the infinite repetitions of the experiment, the state of the composite system $\mathcal{S}+\mathcal{F}$ is $|\Psi^{\mathcal{S}+\mathcal{F}}[m]\rangle$ and the observers $\mathcal{W}$ and $\mathcal{D}$ record the values $l$ and $m$, respectively, immediately after the measurement $\mathcal{M}^{\mathcal{D}}$. We can see that the this situation coincides with that of both Case 1 and Case 2 in Section 13 regarding the measurements $\mathcal{M}^{\mathcal{W}}$ and $\mathcal{M}^{\mathcal{D}}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$, although the statements of both Case 1 and Case 2 in Section 13 are at least operationally vague since they are described in terms of the conventional quantum mechanics where the operational characterization of the notion of probability is not given.
Let us determine the states of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. According to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{W}$ are unchanged after the measurement by the observer $\mathcal{D}$ and therefore the final states of the observer $\mathcal{W}$ are confirmed before the measurement by the observer $\mathcal{D}$, where “apparatus” in Definition 7.3 should read “observer”. Using (101), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{D}}$ is $|\Phi^{\mathcal{W}}[\beta(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}+\mathcal{D}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is